Solve for the net force experienced by q2. (Show your Solutions)

 

 

  q1.    20 µC


  

 


          0.3 m

 

                             

  q2    -15 µC       

                

 

         0.25 m

 

 

 q3 .  -10 µC


1
Expert's answer
2021-02-19T10:32:03-0500

According to the Coloumb's law, the force experienced by "q_2" from "q_1" is:



"F_{12} = k\dfrac{|q_1||q_2|}{r_{12}^2}"

where "k = 9\times 10^9 \space N\cdot m^2/C^2" is the Coloumb's constant, "r_{12} = 0.3m" is the distance between "q_1 = 20\times 10^{-6}C" and "q_2 = -15\times10^{-6}C". Thus, obtain:



"F_{12} = 9\times 10^9\cdot \dfrac{20\times 10^{-6}\cdot 15\times 10^{-6}}{0.3^2} =30N"

Similarly, the the force experienced by "q_2" from "q_3" is:



"F_{12} = k\dfrac{|q_2||q_3|}{r_{23}^2}"

where "r_{23} = 0.25m" is the distance between "q_2" and "q_3 = -10\times10^{-6}C". Thus, obtain:



"F_{13} = 9\times 10^9\cdot \dfrac{20\times 10^{-6}\cdot 10\times 10^{-6}}{0.25^2} =28.8N"

Since "q_1" has the opposite sign with "q_2" and "q_3" have the same sign with "q_2" , both forces are directed in the same direction (toward "q_1"). Thus, according to the superposition principle, the net force will be:



"F_1 = F_{12} + F_{23} = 30N+28.8N = 58.8N"

Answer. 58.8 N.